# OWT: A Foundational Organ-Wise Tokenization Framework for Medical Imaging

### v1

@@ 3.4 Adaptive Holistic Embedding Restorer
Although $\tilde{N}_{g}$ varies per input, this operation preserves organ-specific features while ensuring $X^{\prime}_{H}$ remains a fixed $(dhw)\times c_{e}$ shape, thereby maintaining consistency with the decoder. Finally, we pass $X^{\prime}_{H}$ through a standard decoder [27] ($\mathbf{Dec}$) and unpatchify the output, reconstructing the image $\hat{I}_{T}=\mathbf{Dec}(X^{\prime}_{H})\in\mathbb{R}^{D\times H\times W\times C}$.
As mentioned in Sec. 3.3, our objective is to reduce the discrepancy between the target image $I_{target}$ and the reconstructed image $\hat{I}_{T}$. We combine an L2 term with the LPIPS perceptual loss [62]
@@
4 Experiments
4.1 Datasets and Experimental Setup
@@
Datasets. We conduct experiments on four 3D medical datasets: two CT sets (Abdomen1k [40] and the CT volumes from AutoPet [21]) and two synthetic MRI sequences (delay and pre-artery) from RAOS dataset [39]. These multimodal and multi-sequence datasets validate the effectiveness and robustness of the OWT framework in medical imaging. To further validate that OWT can generalize to different semantic groupings beyond medical scenarios, we employ the CelebAMaskHQ dataset [34]. More details on the preprocessing steps are in Appendix 6.1.
Experimental Setup. All experiments are conducted on six Nvidia H100 GPUs. The hyperparameters, optimizer and scheduler settings follow MAE [27]. We compare OWT against a baseline MAE due to the similar encoder-decoder structure. In OWT, the default token number for each token group is 20. For processing 3D volumes (4 slices by default), we adopt a 3D version of MAE [19]. For fairness, we add LPIPS loss [62] to MAE to improve reconstruction quality. We also explore replacing the encoder-decoder pair with that of VQGAN [17]. Further details of architectures and hyperparameters are in Appendices 6.2 and 6.3, and ablation studies of default settings are in Appendix 8.
4.2 Results
Our OWT framework not only yields highly condensed representations containing rich information but also provides semantically meaningful and disentangled token groups. In this section, we design extensive experiments to evaluate OWT’s performance on several downstream tasks, including reconstruction, segmentation, generalization, and retrieval (Sec. 4.2.1). Notably, our findings indicate that, in contrast to conventional holistic-based methods, OWT can focus on semantic-level tasks and analyses, delivering superior interpretability and controllability. Next, we investigate the characteristics of these semantically disentangled token groups and explore their potential to incorporate other modalities (Sec. 4.2.2). We then demonstrate that the OWT framework can be employed in real-world applications (Sec. 4.2.3). Additionally, we provide ablation studies and explore more architectures of the OWT framework in Appendix 8.
4.2.1 Downstream Analysis
Image Reconstruction. We evaluate the reconstruction capability of OWT across four datasets (two CT and two MRI) [40, 21, 39]. Tables 1 and 2 show that OWT outperforms MAE on L2, LPIPS, and 3D SSIM metrics in holistic-based reconstruction tasks. Moreover, leveraging a patch-embedding module that tokenizes four slices for 3D volumes generates better results than the 2D OWT variant.
@@ 4.2.1 Downstream Analysis
We further explore a semantic-based reconstruction scenario by retaining only a single organ-wise token group at a time, feeding it through AHER and the decoder to generate an organ-specific image. As reported in Tables 1 and 2, using a larger number of slices in CT and MRI volumes leads to more robust disentangled token groups, thereby improving semantic-based reconstruction quality. Fig. 4 illustrates MRI cases (rich in textural detail). Even with LPIPS added, MAE suffers from noticeable detail loss, whereas OWT (all token groups retained) produces sharper and clearer details. Additionally, we can freely combine the token groups to reconstruct only the content of interest.
@@
In the zoom-in rows of Fig. 4, the liver/kidney reconstruction appears nearly identical whether we use only the liver/kidney token group or combine it with others, suggesting that OWT’s organ token groups are highly semantically independent. This property offers better interpretability as well as controllability, enabling reconstruction without unnecessary entanglement of unrelated organ features.
Figure 5: Semantic-based retrieval of CT/MRI images.
Methods CT MRI
@@
UNetplus [63] 88.38 89.29 84.11 84.42
AttU_Net [42] 89.34 89.73 84.54 85.04
TransUnet [9] 88.91 89.70 84.97 84.29
Swin-Unet [5] 86.56 88.06 83.07 82.41
CMUNet [51] 88.33 89.48 84.82 85.56
OWT 89.30 89.52 85.85 85.67
OWT/${}_{\text{4slices}}$ 91.37 +2.03 91.41 +1.68 87.72 +1.87 87.31 +1.64
Table 3: Performance of semantic segmentation.
Image Segmentation. We observe that reconstructing token groups with OWT preserves clear spatial details and boundary information in Fig. 4. Consequently, a simple thresholding operation on the predictions can generate semantic binary masks, offering direct evidence of each organ-wise token group’s independence and interpretability. Table 3 reports the single-organ segmentation Dice scores (averaged over the liver, kidneys, spleen, and pancreas). Overall, the accuracy of OWT’s organ-specific reconstructions is comparable to that of several dedicated 2D segmentation networks. Moreover, incorporating information from adjacent slices (OWT/${}_{\text{4slices}}$) further improves performance, surpassing the second-best approach by an average margin of 1.805 across four CT and MRI datasets (green number in Table 3). The detailed Dice of each organ are in Appendix 7.
AutoPet Holistic-based Rec. 1 Token Group Rec. 1 Token Group Seg.
L2 $\downarrow$ SSIM-3D $\uparrow$ L2 $\downarrow$ SSIM-3D $\uparrow$ Dice $\uparrow$
MAE/${}_{\text{LPIPS}}$ [27] 3.58e-04 0.9872 - - -
AttU_Net [42] - - - - 87.03
CMUNet [51] - - - - 87.89
OWT 1.12e-04 0.9909 4.07e-04 0.9765 87.54
OWT/${}_{\text{4slices}}$ 9.34e-05 0.9924 3.60e-04 0.9798 89.42
@@
Generalization. Table 4 reports the generalization performance when all methods are trained on Abdomen1k and tested on the AutoPet dataset. Under both reconstruction and segmentation metrics, OWT/${}_{\text{4slices}}$ outperforms dedicated holistic-based approaches. Moreover, it also achieves strong results in single-organ reconstruction, indicating that token groups can be flexibly combined or used individually to mitigate interference from irrelevant organs and background regions. Consequently, the semantic features extracted by OWT generalize more effectively, underscoring its potential as a foundational tokenization framework.
@@
4.2.2 Look Deeper into Token Groups
Semantic Information in Token Groups. Fig. 6 illustrates the t-SNE results of how the Organ Collector (with TGR) embeds the background and four organs into the latent space (i.e., five token groups) for each dataset. Within this space, token vectors of the same group type exhibit dense clusters, whereas different token groups remain distinctly separated. This clear inter-group separation highlights the semantic independence of organ-specific token groups in OWT, indicating strong potential for downstream tasks.
Due to the semantic independence of the token group, Fig. 5 shows that our approach not only supports holistic-level retrieval, but also enables organ-level retrieval ($L_{2}$ distance measure). Specifically, we demonstrate top-1, top-2, and maximum-distance retrieval results on both CT and MRI modalities for different token groups. The case indices in the top-left corner of each row indicate how retrieval results vary for different organs of interest. Since experiments in reconstruction and segmentation demonstrate that each token group preserves both the spatial structure and texture details of its corresponding organ, this organ-specific retrieval can be highly beneficial in medical scenarios. For example, it can help radiologists quickly locate similar organ regions across multiple patients, facilitate region-based comparisons of pathology progression, or assist in reference image selection for organ-specific diagnoses.
AutoPet Holistic-based Rec. 1 Token Group Rec.
@@ 4.2.2 Look Deeper into Token Groups
Multi-Modality Incorporation. In OWT, each organ is embedded into its corresponding SDTG, resulting in highly semantically disentangled and independent representations. Therefore, these representations can be efficiently aligned with or collaborated on with other modality embeddings (e.g., text encoding). Specifically, during OWT training, we use medical prompts and a pretrained CLIP encoder [45] to obtain text encoding of each organ, which are then added element-wise to their matching token groups after the Organ Collector. (More details on prompts and architecture for training OWT w/ text encoding are in Appendix 6.2.2.)
Fig. 7 and Table 5 reveal that simply adding text prompts to organ-wise token groups instantly accelerates loss convergence and provides slight improvements in both holistic- and semantic-based reconstruction. Consequently, OWT shows promise for incorporating the text modality to further enhance organ-specific representations. Such multi-modal integration could unlock new medical applications, such as more robust organ labeling, guided retrieval and generation, and improved interpretability in complex workflows.
4.2.3 Organ-wise Token Groups are more than Organs
CelebAMaskHQ Holistic-based Rec. 1 Token Group Rec.
@@ 4.2.3 Organ-wise Token Groups are more than Organs
Figure 8: Samples of holistic-based and semantic-based human face reconstruction on CelebAMaskHQ.
Figure 9: Semantic-based retrieval of CelebAMaskHQ.
@@ 5 Conclusion
In this paper, we propose a novel Organ-Wise Tokenization (OWT) framework and Token Group-Based Reconstruction (TGR) paradigm, which disentangle holistic embeddings into semantically disentangled token groups. Our approach not only improves interpretability, generalizability, efficiency and controllability in organ-based analyses, but also shows adaptability to broader scenarios, including natural image domains. By enabling semantic-level token groups, OWT enriches downstream tasks such as semantic-based reconstruction, segmentation and retrieval, offering a new perspective on how representations can be leveraged.
@@
Abdomen1k Liver Kidney Spleen Pancreas Avg Dice $\uparrow$
UNetplus [63] 96.85 94.63 91.00 71.05 88.38
AttU_Net [42] 97.09 95.22 93.05 72.00 89.34
TransUnet [9] 96.78 94.65 92.30 71.89 88.91
SwinUnet [5] 96.53 93.68 90.52 65.50 86.56
CMUNet [51] 96.05 94.67 90.84 71.75 88.33

### v2

@@ 2.4 Adaptive Holistic Embedding Restorer
Although $\tilde{N}_{g}$ varies per input, this operation preserves organ-specific features while ensuring $X^{\prime}_{H}$ remains a fixed $(dhw)\times c_{e}$ shape, thereby maintaining consistency with the decoder. Finally, we pass $X^{\prime}_{H}$ through a standard decoder [he2022masked] ($\mathbf{Dec}$) and unpatchify the output, reconstructing the image $\hat{I}_{T}=\mathbf{Dec}(X^{\prime}_{H})\in\mathbb{R}^{D\times H\times W\times C}$.
As mentioned in Sec. 2.3, our objective is to reduce the discrepancy between the target image $I_{target}$ and the reconstructed image $\hat{I}_{T}$. We combine an L2 term with the LPIPS perceptual loss [zhang2018unreasonable]
@@
Tasks Organ-specific Organ-level Pre-defined Any Comb. Segment-
Identification Retrieval Rec. Rec. ation
Rec. Nets
Seg. Nets
[…]
Table 1: Downstream task capabilities after one-time training.
3 Datasets and Experimental Setup
@@
We conduct experiments on four 3D medical datasets: two CT sets (Abdomen1k [ma2021abdomenct] and the CT volumes from AutoPet [Gatidis2022]) and two synthetic MRI sequences (delay and pre-artery) from RAOS dataset [luo2024rethinking]. These multimodal and multi-sequence datasets validate the effectiveness and robustness of the OWT framework in medical imaging. To further validate that OWT can generalize to different semantic units beyond medical scenarios, we employ the CelebAMaskHQ dataset [lee2020maskgan]. More details are in Appendix B.1, including image preprocessing and data splitting.
All experiments are conducted on six Nvidia H100 GPUs. For fairness, we also add LPIPS loss [zhang2018unreasonable] to VAE and MAE to improve the quality of reconstruction. Further details of the architecture descriptions and hyperparameters are in Appendix B.2 and B.3.
Methods Enc. OC. TG Enc. AHER Dec. Total
MAE 16.66 - - - 11.11 27.77
[…]
Table 2: Comparison of computation costs (GFLOPs).
Methods Disentangled Used Tumor Identification
Organ Info. Token Num. Liver Kidneys
MAE 196 73.5 74.7
[…]
Table 3: Performance for the tumor identification tasks.
4 Results
In this section, we design extensive experiments to explore OWT’s unique clinical impacts (Sec. 4.2), and to evaluate performance on several downstream tasks (Sec. 4.3). Next, we dive into SDTGs and investigate their semantic characteristics (Sec. 7). We also explore the potential of our OWT to address the common issue of limited medical annotations (Sec. 4.5). Finally, we demonstrate that the OWT framework can also be employed in natural image applications (Sec.4.6), and provide ablation studies in Appendix D to discuss several important hyperparameters of OWT, including encoder/decoder alterations, different token number per organ, and data percentage used for training.
Abdomen1k Holistic Rec. 1 Token Group Rec.
[…]
MAE 6.74e-04 0.0684 0.9688
[…]
AutoPet Holistic Rec. 1 Token Group Rec.
[…]
MAE 2.86e-04 0.0452 0.9799
[…]
Table 4: Performance of image reconstruction on CT datasets.
Delay Holistic Rec. 1 Token Group Rec.
[…]
MAE 1.37e-03 0.0465 0.9686
[…]
PreArtery Holistic Rec. 1 Token Group Rec.
[…]
MAE 1.67e-03 0.0438 0.9701
[…]
Table 5: Performance of image reconstruction on MRI datasets.
4.1 Advances in Design and Architecture
Due to our proposed TGR training paradigm, the OWT framework produces semantically disentangled token groups (SDTGs) rich in organ-specific information. Leveraging the AHER module, these SDTGs can be individually or flexibly combined and restored from semantic embeddings into their original spatial dimensions. Consequently, as shown in Table 1, OWT requires only one-time training to enable multiple downstream tasks. This cross-task efficiency and controllability significantly distinguish OWT from conventional networks with entangled embeddings, which cannot achieve such flexibility.
OWT is also computationally efficient. As shown in Table 2, OWT requires only 24.01, notably fewer than MAE’s 27.77 GFLOPs. This efficiency stems from (1) OWT’s semantic tokens containing more highly condensed information than the traditional entangled holistic tokens, and (2) the lightweight design of the Organ Collector (0.25 GFLOPs) and AHER (0.074 GFLOPs) modules.
These designs also offer substantial potential. Specifically, we explore adaptively distributing different token numbers to SDTGs (e.g., based on organ volumes) in Appendix D.3. Since Organ Collector and AHER are encoder-decoder independent, we further investigate notable performance gains by replacing Transformer-based structures with VQGAN-based alternatives in Appendix D.1.
Figure 5: Semantic-based organ-level retrieval of MRI images.
4.2 Unique Clinical Impacts
Organ-Specific Tumor Identification. In contrast to conventional models with entangled holistic embeddings, OWT extracts organ-specific information into distinct, semantically disentangled token groups (SDTGs). This structure enables the efficient use of only relevant tokens for highly specific downstream tasks. We demonstrate this on a classification task derived from tumor labels in the Abdomen1k dataset. As shown in Table 3, OWT leverages only 20 organ-specific tokens (i.e., for liver or kidneys) yet achieves significantly better performance than the holistic embedding-based MAE, which utilizes 196 tokens. This results in accuracy improvements of 6.2% (79.7 vs. 73.5) and 7.1% (81.8 vs. 74.7) for identifying tumors in the liver and kidneys, respectively, highlighting OWT’s ability to create condensed yet highly informative representations for targeted clinical tasks.
Organ-Level Retrieval. Due to their semantic independence, our token groups enable organ-level retrieval ($L_{2}$ distance measure), a capability holistic embedding-based models cannot perform (Fig. 5). We demonstrate top-1, top-2, and maximum-distance retrieval results on the MRI modality (CT examples are in Appendix F), where case indices in the top-left corner show results for different organs of interest. Since each token group preserves its organ’s spatial structure and texture details, this targeted retrieval is highly beneficial in medical scenarios. For example, it can help radiologists quickly locate similar organ regions across patients without being influenced by other organs, facilitate region-based pathology comparisons, or assist in selecting reference images for organ-specific diagnoses.
4.3 Downstream Analysis
[…]
@@ 4.3 Downstream Analysis
We further explore semantic-based reconstruction by retaining a single organ-wise token group, which is fed through AHER and the decoder to generate an organ-specific image. As reported in Tables 4 and 5, using more slices leads to more robust token groups and improves semantic reconstruction quality. Fig. 4 illustrates this on MRI cases: even with LPIPS added, MAE suffers from noticeable detail loss (red arrows), whereas OWT (all token groups retained) produces sharper, clearer details. Additionally, token groups can be freely combined to reconstruct only content of interest.
@@
The zoom-in rows of Fig. 4 show that the liver/kidney reconstruction appears nearly identical whether using only its specific token group or combining it with others, suggesting OWT’s organ token groups are highly semantically independent. This property enhances interpretability and controllability, enabling reconstruction without unnecessary entanglement of unrelated organ features.
1- / 4-slice CT MRI
@@
UNetplus 88.38 / 89.78 89.29 / 91.04 84.11 / 85.50 84.42 / 85.68
AttU_Net 89.34 / 90.75 89.73 / 91.46 84.54 / 86.75 85.04 / 86.52
TransUnet 88.91 / 90.24 89.70 / 91.34 84.97 / 86.40 84.29 / 86.01
Swin-Unet 86.56 / 89.47 88.06 / 89.91 83.07 / 85.39 82.41 / 84.48
CMUNet 88.33 / 90.92 89.48 / 91.50 84.82 / 86.82 85.56 / 86.94
[…]
Table 6: Dice performance (1- / 4-slice) of semantic segmentation.
AutoPet Holistic Rec. 1 TG Rec. 1 TG Seg.
[…]
AttU_Net 87.03
CMUNet 87.89
[…]
@@
Image Segmentation. Reconstructing OWT’s token groups preserves clear spatial details and boundaries (Fig. 4). Consequently, a simple thresholding operation on the predictions can generate semantic binary masks, offering direct evidence of each organ-wise token group’s independence and interpretability (More details on the thresholding operation are in Appendix B.3). Table 6 reports the single-organ segmentation Dice scores (averaged over the liver, kidneys, spleen, and pancreas). For fair comparison, all networks report segmentation results based on 1 and 4 slices. Overall, the accuracy of OWT incorporating information from adjacent slices (OWT with 4-slice) achieves comparable or superior performance across CT and MRI datasets. Detailed per-organ Dice comparisons are in Appendix E.
[…]
4.4 Semantic Information in Token Groups
@@
AutoPet Holistic Rec. 1 TG Rec.
@@
T-SNE Analysis. Fig. 6 illustrates the t-SNE results of how the Organ Collector (with TGR) embeds the background and four organs into the latent space (i.e., five Token Groups) for each dataset. Within this space, token vectors of the same group type exhibit dense clusters, whereas different token groups remain distinctly separated. This clear inter-group separation highlights the semantic independence of organ-specific token groups in OWT.
Multi-Modality Incorporation. OWT’s highly disentangled and independent SDTGs enable efficient alignment or incorporation with other modalities, such as text. We demonstrate this by encoding medical prompts with a CT-CLIP encoder (specifically pretrained on paired CT images and reports) [hamamci2024developing], and adding the resulting organ-specific and CT-consistent text embeddings element-wise to their corresponding token groups. More details on prompts and architecture for training OWT with text encoding are in Appendix C.
As shown in Fig. 7 and Table 8, this simple integration accelerates loss convergence and provides immediate improvements in several downstream tasks. This highlights OWT’s promise for multi-modal enhancement, unlocking potential applications, such as text-guided retrieval, generation, and improved interpretability in complex workflows.
AutoPet Data Holistic Rec. 1 TG Rec. Seg.
[…]
OWT 100% 1.49e-04 0.0232 3.32e-04 0.0071 0.8952
OWT 20% 5.54e-04 0.0658 5.70e-04 0.0115 0.8109
[…]
OWT 60% 1.86e-04 0.0270 3.49e-04 0.0076 0.8858
[…]
Table 9: Performance comparison of OWT in different training strategies.
CelebAMaskHQ Holistic Rec. 1 Token Group Rec.
@@
Figure 8: Samples of holistic and semantic-based human face reconstruction on CelebAMaskHQ.
4.5 Semi-Supervised Training under Limited Annotations
The limitation of annotation is a long-standing issue in medical imaging. In our foundational framework, we find that appropriate supervision of token groups is beneficial for downstream tasks. Therefore, we adopted a two-step training strategy: (1) using all data without labels for unsupervised pretraining of OWT’s encoder and decoder via whole image reconstruction for 600 epochs, and (2) fine-tuning the entire OWT with only partial annotation coverage (20% and 60%) from the training set for 1200 epochs.
As shown in Table 9, this approach significantly improves performance under label scarcity. The semi-supervised model using 20% labels greatly outperformed the model trained on only 20% labels. Furthermore, the semi-supervised mode with 60% labels nearly matched the upper-bound results of training with 100% labels , showing OWT’s robustness and potential for data-scarce scenarios.
4.6 Token Group as Compositional Units
Figure 9: Semantic-based retrieval of CelebAMaskHQ.
@@ 5 Conclusion
In this paper, we propose a novel Organ-Wise Tokenization (OWT) framework and Token Group-Based Reconstruction (TGR) paradigm, which disentangle holistic embeddings into semantically disentangled token groups. Our approach not only improves interpretability, generalizability, efficiency and controllability in organ-based analyses, but also shows adaptability to broader scenarios, including natural image domains. By enabling semantic-level token groups, OWT enriches unique clinical applications and downstream tasks such as organ-specific analysis, semantic-based reconstruction, segmentation and retrieval, offering a new perspective on how representations can be utilized.
[…]
@@
Appendix A Related Work
A.1 Holistic Embedding-based Representation Learning
Holistic embedding-based approaches map entire 2D/3D inputs into a single latent representation, aiming to capture global image features [voulodimos2018deep, geirhos2020shortcut, wang2024disentangled]. These methods typically follow either self-supervised or task-guided strategies. In self-supervised learning, contrastive techniques like SimCLR [chen2020simple], SimSiam [chen2021exploring], and BYOL [grill2020bootstrap] enforce consistency between augmented views, while predictive approaches such as CPC [oord2018representation] and RotNet [gidaris2018unsupervised] reconstruct missing or permuted portions. Recently, masked image modeling (e.g., MAE [he2022masked] and BEiT [bao2021beit]) has emerged as an effective approach by requiring models to infer masked regions. Meanwhile, task-guided frameworks focus on extracting features that directly benefit a target task, typically resulting in holistic embeddings. Many classic neural networks, such as ResNet [he2016deep], DenseNet [huang2017densely], and Vision Transformer [dosovitskiy2020image], and their variants incorporate classification, regression, or segmentation objectives.
In medical imaging, holistic representations have been widely used for self-supervised feature learning in CT [chen2021momentum, lai2024e3d], MRI [chaitanya2020contrastive, lang20233d], and X-ray [tiu2022expert, xiao2023delving]. Additionally, classical supervised pipelines (e.g., U-Net [ronneberger2015u] and its variants [zhou2018unet++, chen2021transunet, cao2022swin, hatamizadeh2021swin]) rely on a global bottleneck to embed entire images or volumes, demonstrating strong performance in organ segmentation and disease detection. Although such approaches achieve notable success, they often encode multiple semantic factors, such as multiple organs and the surrounding background, into an entangled representation with limited interpretability and fine-grained control [schutte2021using, cetin2023attri, wang2024disentangled].
A.2 Disentangled Representation Learning
In contrast to holistic embedding-based approaches, various disentangled representation learning (DRL) studies have pursued a dimension-wise strategy for two main objectives. First, to derive more robust holistic embeddings with improved generalizability [dangovski2021equivariant, bai2023robust, devillers2023equimod, wang2024distortion, garrido2023self, song2024contrastive, wang2024disentangled], and second, to explicitly decompose images into latent factors, each corresponding to distinct and semantically meaningful attributes, in order to improve explainability [burgess2018understanding, chen2016infogan, xiao2017dna, zhu2018visual].
In medical imaging, the motivations for DRL are especially strong: improved interpretability, controllability, and robustness to confounding factors are crucial in clinical settings [liu2022learning, wang2022disentangled]. When transitioning DRL to medical imaging, researchers seek to decouple anatomical properties (e.g., organ size or shape) [wang2022semantic, jiang2022disentangled], modality characteristics (e.g., CT and MRI sequences) [ouyang2021representation, chartsias2019disentangled], and pathological factors (e.g., tumor presence or severity) [liu2022joint, hu2022clinically], aiming to isolate distinct clinical features.
Despite notable progress, data- and perturbation-driven DRL approaches can be sensitive to training objectives, architecture choices, and assumptions about independent factors [locatello2019challenging, wang2024distortion, wang2024disentangled]. In existing studies, disentangled dimensions may not align precisely with real anatomical or pathological concepts, and spatial correspondence often remains weak when images are globally encoded. Consequently, there has been a growing emphasis on methods that combine disentanglement principles with controlled grouping or tokenization schemes to encapsulate semantically meaningful subsets of features.
Appendix B Implementation Details
B.1 Image Preprocessing
For all 3D medical volumes, we standardize the spatial resolution to $2.0\text{mm}\times 2.0\text{mm}\times 2.0\text{mm}$ and apply a center crop of $224\times 224$. For CT datasets, we clip intensity values to the [1000,1000][-1000,1000] Hounsfield Unit (HU) range and normalize them to $[0,1]$. For MRI datasets, we clip intensities to the $[0,99]$ percentile range and normalize them to $[0,1]$. We define five token groups to extract semantic features based on common annotations: background, liver, kidneys, spleen, and pancreas. Each medical image is further standardized in thickness by center-padding or cropping 112 slices based on the anatomical centers of the four organs.
For the Abdomen1K dataset [ma2021abdomenct], the number of cases in the training and test sets is 899:101. For AutoPet [Gatidis2022], the split is 700:200, while for RAOS delay and pre-artery sequences [luo2024rethinking], both have a training-to-test ratio of 330:83. All experiments, including OWT, MAE, and segmentation networks, are based on the same preprocessed dataset.
For the 2D face images in CelebAMaskHQ dataset [lee2020maskgan], we resize them to $224\times 224$, and define six semantic entities based on annotations (background, face&ear, hair&hat, mouth, nose, and eyes&glasses), as shown in Fig. 8. The CelebAMaskHQ dataset contains 30,000 facial image cases, with 24,000 images allocated for training and 6,000 for testing.
B.2 Default Architecture of OWT and MAE
In OWT, we adopt the same encoder and decoder architecture as MAE [he2022masked], with the only modification being the replacement of the self-attention blocks [vaswani2017attention] with linear self-attention blocks [katharopoulos2020transformers]. The hyperparameters, optimizer and scheduler settings all follow MAE. By default, we use the ViT-base architecture, where the MAE encoder consists of 12 attention blocks, the decoder has 8, and the token embedding dimension is set to 768. For OWT, we configure the Encoder and Token Group Encoder with 6 attention blocks each, while keeping the decoder at 8 blocks. This ensures that the overall model length remains consistent with MAE.
The default token number for each token group is 20 in OWT. Therefore, the training load of OWT is lower than that of MAE when using fewer than 10 SDTGs. To prevent information leakage during TGR training, we remove the class token from the OWT architecture. For OWT processing 2D images, we use a $16\times 16$ kernel for patch embedding. For 3D input volumes, we apply a $1\times 16\times 16$ 3D kernel for patch embedding.
Method AutoPet CelebAMaskHQ
L2 $\downarrow$ LPIPS $\downarrow$ L2 $\downarrow$ LPIPS $\downarrow$